The density of NO2 in a 4.77 L tank at 738.4 torr and 45.0 °C is ________ g/L.
Answer #1

P= 738.4 torr

= (738.4/760) atm

= 0.9716 atm

T= 45.0 oC

= (45.0+273) K

= 318 K

Molar mass of NO2,

MM = 1*MM(N) + 2*MM(O)

= 1*14.01 + 2*16.0

= 46.01 g/mol

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

0.9716 atm *46.01 g/mol = density * 0.08206 atm.L/mol.K *318.0 K

density = 1.71 g/L

Answer: 1.71 g/L